tell whether the system has one solution, infinitely many solutions,or no solution? x+y+3=0
3x-2y+4+0

Answer by LinnW(1048)

x+y+3=0
3x-2y+4=0
x + y = -3
3x - 2y = -4
multiply the top equation by 2
2x + 2y = -6
3x - 2y = -4 add these together
5x = -10
x = -2
Now substitute in
x + y = -3
-2 + y = -3
add 2 to each side
y = -1
So x = -2 and y = -1